16. Calculate What is the volume of an ideal gas if there are 4.10 X 102 particles at a temperature of 262 K and a pressure of 195 kPa?




1
Expert's answer
2021-05-12T08:38:03-0400

According to the key equation in kinetic molecular theory, the pressure is


"p=\frac NVkT,\\\space\\
V=\frac{NkT}{p}=\frac{(4.1·10^2)(1.38·10^{-23})262}{195·10^3 }=7.61·10^{-24}\text{ m}^3."